OpenStudy (anonymous):

Which ratios are proportional to 27\30

OpenStudy (anonymous):

canceling 3 we get 9/10

OpenStudy (texaschic101):

ratios are nothing but equivalent fractions 27/30 reduces to 9/10....so the ratio 9:10 is proportional or 27/30 * 2/2 = 54/60....the ratio 54:60 is proportional 27/30 * 3/3 = 81/90...the ratio 81:90 is proportional